OpenStudy (anonymous):

Solve 11.32x - 1 = 15.7 (2 points) Answer x ≈ 1.12 x ≈ 0.57 x ≈ 0.94 x ≈ 1.07

OpenStudy (anonymous):

1. First you wanna isolate x onto one side on it's own or in other words " group like terms together" 11.32x = 15.7 +1 11.32x = 16.7 2.Divide both sides by 11.32 in order to isolate x completely by itself x = 16.7/11.32 x should be around 1.48...

OpenStudy (anonymous):

That's not an option though :(